Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost.

Last year, the company sold 44,000 of these balls, with the following results:

Sales (44,000 balls)

$

1,100,000

Variable expenses

660,000

Contribution margin

440,000

Fixed expenses

317,000

Net operating income

$

123,000

Required:

  1. Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level.

Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level. (Round "Unit sales to break even" to the nearest whole unit and other answers to 2 decimal places.)

CM Ratio

%

Unit sales to break even

balls

Degree of operating leverage

  • Req 1
  • Req 2
  1. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

Ratio" to 2 decimal places and "Unit sales to break even" to the nearest whole unit.)

CM Ratio

%

Unit sales to break even

balls

  • Req 1
  • Req 3
  1. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $123,000, as last year?

Number of balls

  • Req 2
  • Req 4
  1. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

Req 3

Selling price

  • Req 3
  • Req 5
  1. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

CM Ratio

%

Unit sales to break even

balls

  • Req 4
  • Req 6A

6. Refer to the data in (5) above.

a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $123,000, as last year?

Question: What is the number of balls?

b. Assume the new plant is built and that next year the company manufactures and sells 44,000 balls (the same number as sold last year). Prepare a contribution format income statement and compute the degree of operating leverage.

Northwood Company

Contribution Income Statement

Degree of operating leverage

  • Req 6A_QC_CS-184684

Brewer 8e Rechecks 2018-09-04

Brewer 8e Rechecks 2019-08-29
Answer #1
Answer
1 Contribution Margin = Sales - Variable Cost / Sales *100
Particulars Basketballs
Sales Qty 44000
Sales price per unit $                    25.00
Sales Value $      1,100,000.00
Less: Variable Cost
Variable cost per unit $                    15.00
Variable Cost $          660,000.00
Contribution = Sales - Variable Cost $          440,000.00
Contribution Margin = 440,000/1,100,000*100
Contribution Margin = 40%
Contribution $          440,000.00
Less: Fixed Cost $          317,000.00
Net Operating Profit $          123,000.00
Unit Sales to Break even
in Units = Fixed Cost / Contribution per unit
Contribution per unit=
Sales price per unit $                    25.00
Variable cost per unit $                    15.00
Contribution per unit= $                    10.00
in Units = Fixed Cost / Contribution per unit 317000/10
Unit Sales to Break even =                31,700.00
Degree of operating leverage = Contribution / EBIT
Contribution =              440,000.00
EBIT=              123,000.00
Degree of operating leverage =                           3.58
2 Sales Qty 44000
Sales price per unit $                    25.00
Sales Value $      1,100,000.00
Less: Variable Cost
Variable cost per unit $                    18.00
Variable Cost $          792,000.00
Contribution = Sales - Variable Cost $          308,000.00
Contribution Margin = 308,000/1,100,000*100
Contribution Margin = 28%
Contribution $          308,000.00
Less: Fixed Cost $          317,000.00
Net Operating Profit $            (9,000.00)
Unit Sales to Break even
in Units = Fixed Cost / Contribution per unit
Contribution per unit=
Sales price per unit $                    25.00
Variable cost per unit $                    18.00
Contribution per unit= $                       7.00
in Units = Fixed Cost / Contribution per unit 317000/7
Unit Sales to Break even =                      45,286
3 Refer to the above date 2
Bolls will have to sold next year to earn the same net operating income, $123,000 as last year as follows
Fixed Cost + Profit / contribution per unit
ie =( 317000 + 123000) / 7                      62,857
4 Selling price of $30 must charge next year to cover the increase labor cost
Workings
let X = Selling price X - 18 / X = .40
X = 30
Sales Qty 44000
Sales price per unit $                    30.00
Sales Value $      1,320,000.00
Less: Variable Cost
Variable cost per unit $                    18.00
Variable Cost $          792,000.00
Contribution = Sales - Variable Cost $          528,000.00
Contribution Margin = 528,000/1,320,000*100
Contribution Margin = 40%
5 Sales Qty 44000
Sales price per unit $                    25.00
Sales Value $      1,100,000.00
Less: Variable Cost
Variable cost per unit = 15-40% $                       9.00
Variable Cost $          396,000.00
Contribution = Sales - Variable Cost $          704,000.00
Contribution Margin = 704,000/1,100,000*100
Contribution Margin = 64%
Contribution $          704,000.00
Less: Fixed Cost (Increased to double) $          634,000.00
Net Operating Profit $            70,000.00
Unit Sales to Break even
in Units = Fixed Cost / Contribution per unit
Contribution per unit=
Sales price per unit $                    25.00
Variable cost per unit $                       9.00
Contribution per unit= $                    16.00
in Units = Fixed Cost / Contribution per unit 634000/10
Unit Sales to Break even =                39,625.00
6.a) Bolls will have to sold next year to earn the same net operating income, $123,000 as last year as follows
Fixed Cost + Profit / contribution per unit
ie =( 634,000 + 123000) / 16                      47,313
6.b) Northwood Company
Contribution Income Statement
Particulars Basketballs
Sales Qty 44000
Sales price per unit $                    25.00
Sales Value $      1,100,000.00
Less: Variable Cost
Variable cost per unit $                       9.00
Variable Cost $          396,000.00
Contribution = Sales - Variable Cost $          704,000.00
Less: Fixed Cost $          634,000.00
Net Operating Profit $            70,000.00
Degree of operating leverage = Contribution / EBIT
Contribution =              704,000.00
EBIT=                70,000.00
Degree of operating leverage =                         10.06